# Neural Guided Diffusion Bridges

### v2

@@ B.4 FitzHugh-Nagumo model
Normal event: We set $\delta t=0.005$, which leads to $N=400$ time steps. The observation noise variance is $\epsilon=1e-4$ to ensure a stable backward filtering ODE solution. For the neural guided bridge, $\vartheta_{\theta}(t,x)$ is constructed as a fully connected neural network with 4 hidden layers and 32 hidden dimensions at each layer, activated by tanh functions. The training is done with 225,000 independent samples from $X^{\bullet}_{t}$, batch size $M=30$, optimized by Adam with an initial learning rate of $45e-$ 4 and a cosine decay scheduler. For the guided proposal, we set $\eta=0$ as suggested in Bierkens et al. [2020] and run one chain for 50,000 iterations with a burn-in of 20,000 steps, obtaining $61.15\%$ acceptance rate. The reference is obtained by brutal-forced sampling $X$ 100,000 times and applying the filtering, which obtain 4814 valid samples, only the first 30 are shown in the figure.
Rare event: We use the same time step size as the previous experiment, however, a quadratic transformation $\tau(t)=t(2-t/T)$ is deployed to get an irregular grid, where the closer to the endpoint $T$, the denser the grid. The observation noise variance is $\epsilon=2e-4$. $\vartheta_{\theta}(t,x)$ architecture is the same, but training for a longer time with 300,000 independent samples. For the guided proposal, $\eta=0.9$, it runs one chain for 20,000 iterations with a burn-in period as the first 2,000 steps and obtains $23.18\%$ acceptance rate. Another notable thing is to obtain the reference by brutal forced sampling, one can only get 104 ($1.04$) samples that meet the rare conditionals out of 100,000 forward samplings, while in the normal case, the number of valid samples is 4814 ($4.81\%$).
@@ B.5 Stochastic landmark matching
To demonstrate the necessity of using a state-dependent nonlinear process as Equation (30) to model the shape process rather than a linear one with a fixed diffusion matrix, we show a comparison of forward processes under two scenarios as Figure 13, where the linear process is similar to the auxiliary process Equation (31), but with a fixed $Q$ evaluated at the starting point $x_{0}$, except for the evaluation point, the rest parameters for the kernel are the same. In Equation (30), as $X_{t}^{(j)}$ approaching $X_{t}^{(i)}$, the correlation between them becomes stronger, which forces them to move synchronically, thus preventing the overlapping and intersection. Also, when $\kappa\to 0$, the process defined by Equation (30) will degenerate to Brownian motion.
Figure 13: Comparison of using a linear (Top row) with a fixed diffusion coefficient and Equation (30) (Bottom row) to model the stochastic evolution of an ellipse shape (marked in purple) using 4 different samples of trajectories individually, where the nonlinear Equation (30) preserves the topology of the ellipse, which is broken during the evolution under the linear process.
The observation noise variance is set as $\epsilon=2e-3$, as we find too small values of $\epsilon$ will cause numerical instability. We deploy the neural network architecture suggested in Heng et al. [2022] to model $\vartheta_{\theta}$, whose encoding part is a two-layer MLP with 128 hidden units at each layer, and the decoding part is a three-layer MLP with hidden units of 256, 256, and 128 individually. The network is activated by tanh, and trained with 240,000 independent samples from $X^{\bullet}$ with batch size $N=8$, optimized by Adam with an initial learning rate of 7.0e-4 and a cosine decay scheduler. For the guided proposal, we run one chain for 5,000 iterations and drop the first 1,000 iterations as the burn-in, with $\eta=0.95$ and obtain $12.62\%$ acceptance rate on average.

### v3

@@ B.4 FitzHugh-Nagumo Model
Normal event: We set $\delta t=0.005$, which leads to $M=400$ time steps. We took $\epsilon^{2}=1e-8$. For the neural guided bridge, $\vartheta_{\theta}(t,x)$ is constructed as a fully connected neural network with 4 hidden layers and 32 hidden dimensions at each layer, activated by LipSwish functions. The training is done for $5,000$ gradient descent steps. In each step, we generated $N=100$ independent samples from the current $X^{\bullet}$ for Monte Carlo estimation. The network is optimized by Adam with a learning rate of $31e-$ 3. For the guided proposal, we set $\eta=0$ as suggested in Bierkens et al. [2020] and ran one chain for $50,000$ iterations with a burn-in of $20,000$ steps, obtaining $64.41\%$ acceptance percentage. The reference is obtained by sampling the (unconditional) forward process filtering samples with $\|LX_{T}-v\|\leq 0.01$, as $v=\begin{bmatrix}1.0\end{bmatrix}$ is an event around which the process is likely to reach a small ball, one can expect to easily obtain sufficient samples that meets the filtering condition, it turns out that we only need to sample the forward process for $10,000$ to obtain $450$ vaild samples, and we only show the first $30$ in the figure.
(a) neural bridge
(b) guided proposal
(c) reference
Figure 11: Visualization of 30 simulated bridge trajectories under normal event conditioning from the learned neural bridge and pCN sampling of the guided proposal. The reference trajectories in (c) are obtained by forward sampling the unconditioned process, keeping only samples that satisfy the condition $\|LX_{T}-v\|\leq 0.01$.
Rare event: We duplicate the neural network training setting as before, including the used network structure and training hyperparameters. For the guided proposal we set the pCN-parameter to $\eta=0.9$ and ran one chain for $50,000$ iterations which yielded a Metropolis-Hastings acceptance percentage equal to $22.46\%$. The initial $20,000$ iterations are considered burnin samples. As a reference, we forward sampled $200,000$ paths of the unconditioned process. Of those, $35$ samples satisfied $\|LX_{t}-v\|<0.01$, which is about $0.02\%$ of the samples. As expected, in the “normal event” case, this percentage was higher ($4.5\%$).
(a) neural bridge
(b) guided proposal
(c) reference
Figure 12: Visualization of 30 simulated bridge trajectories under rare event conditioning from the learned neural bridge and pCN sampling of the guided proposal. The reference trajectories in (c) are obtained by forward sampling the unconditioned process, and filtered the results by the condition $\|LX_{T}-v\|\leq 0.01$.
@@ B.5 Stochastic Landmark Matching
Figure 13: Comparison of unconditioned samples from the linear process (Equation 33) and the nonlinear process (Equation 32). Top row: $4$ independent samples from Equation 33. Bottom row: corresponding samples from Equation 32, each using the same Wiener process realization as in the top row.
The observation noise variance is set as $\epsilon^{2}=2e-3$, as we find too small values of $\epsilon$ will cause numerical instability. We deploy the neural network architecture suggested in Heng et al. [2022] to model $\vartheta_{\theta}$, whose encoding part is a two-layer MLP with 128 hidden units at each layer, and the decoding part is a three-layer MLP with hidden units of 256, 256, and 128 individually. The network is activated by tanh, and trained with 240,000 independent samples from $X^{\bullet}$ with batch size $N=8$, optimized by Adam with an initial learning rate of $47.0e-$ 4 and a cosine decay scheduler. For the guided proposal, we run $4$ chain for $5,000$ iterations each and drop the first $1,000$ iterations as the burn-in, with $\eta=0.95$ and obtain $12.62\%$ acceptance percentage on average over the 4 chains.
Figure 14: $4$ realizations of sampling from the guided proposal, the trained neural bridge and the final iteration after updating the guided proposal with $5,000$ pCN-steps. Top row: $4$ samples of the guided proposal using $4$ independent Wiener process realisations; middle row: $4$ samples from the trained neural bridge using the same Wiener realisations as the top row; Bottom row: the final outputs of $4$ independent chains updating the guided propsals using pCN-steps. The chains are initialised with the same Wiener realisations as the top row.